# On a counterexample to Mordell’s Pellian equation conjecture: a non-computer based approach

### v1

@@ Abstract.
In this note, we investigate a recently discovered counterexample to Mordell’s Pellian equation conjecture. We provide a verification of this counterexample that can be checked without computer assistance.
@@ 1. Introduction and main result
Observe that $\mathcal{O}_{K}=\mathbb{Z}[\omega]=\mathbb{Z}\oplus\omega\mathbb{Z}$. Let $p\in\mathbb{P}$. Then $p$ is called ramified in $\mathcal{O}_{K}$ if $p\mathcal{O}_{K}$ is the square of a maximal ideal of $\mathcal{O}_{K}$. Moreover, we say that $p$ splits in $\mathcal{O}_{K}$ if $p\mathcal{O}_{K}$ is the product of two distinct maximal ideals of $\mathcal{O}_{K}$. After introducing the most important terminology above, we are now prepared to formulate Mordell’s Pellian equation conjecture.
Conjecture 1.1.
(Mordell’s Pellian equation conjecture) Let $x,y\in\mathbb{N}_{0}$ be such that $\varepsilon=x+y\omega$. If $d\in\mathbb{P}$ with $d\equiv 3\mod 4$, then $d\nmid y$.
This conjecture was stated by L. J. Mordell in 1961 ([5, page 283]) and independently by A. A. Kiselev and I. Š. Slavutskiĭ in 1959 ([4]). It was investigated in various papers and by many authors (for instance, see [1, 3, 8, 9]). Next, we present the main result of this note.
@@
The counterexample in Theorem 1.2 was first announced in [7] and it was discovered and verified with computer assistance. For the methods, algorithms and programs that were used to find (and check) this counterexample, we refer to [7, Section 3].
The purpose of this note is to provide a detailed proof of Theorem 1.2 that can be checked without relying on computers. That said, the proof in Section 3 was created with computer assistance (and with the help of the programs Mathematica 12.0.0 and Pari/GP 2.15.2). We want to outline why (we think that) this proof can be verified without computers (within a manageable amount of time). Note that the vast majority of the verification process involves doing basic arithmetic. Therefore, we focus entirely on the arithmetic in the estimate below. We define an arithmetical operation as hard if it is a multiplication of integers $a$ and $b$ with $|a|\geq 10^{3}$, $|b|\geq 10^{3}$ and $|ab|\geq 10^{10}$. All other arithmetical operations in this note (additions, subtractions and multiplications of integers) are defined to be easy. We also define additions and subtractions where (at least) one of the members is a single digit integer as trivial. A rough estimate yields that there are approximately 550 hard operations and about 2200 (nontrivial) easy operations in the proof of Theorem 1.2 below. If we assume that (the verification of) a hard operation takes about 15 minutes (on average) and (the verification of) an easy (but nontrivial) operation takes about 3 minutes (on average), then a verification of the proof below is doable in about 250 hours. (That said, we think that skilled arithmeticians can perform the verification of all details below in substantially less than 100 hours.)
@@ 2. Preliminaries
We recall a variant of a result of H. C. Pocklington (cf. [2, Theorem 4] and [6]) that will be used to show that $d$ is prime. For the readers’ convenience we include a proof. This proof goes along the lines of the proof of [2, Theorem 4 and Corollary 1].
@@
Let $d=39028039587479$ and let $x,y\in\mathbb{N}_{0}$ be such that $\varepsilon=x+y\omega$. We need to show that $d\equiv 3\mod 4$, $d\in\mathbb{P}$ and $d\mid y$. Since $79=19\cdot 4+3$, we have that $d\equiv 79\equiv 3\mod 4$. Observe that $\omega=\sqrt{d}$ and $\varepsilon=x+y\sqrt{d}$.
@@
$p=$ $1808\cdot 2+1=1205\cdot 3+2=723\cdot 5+2=516\cdot 7+5=328\cdot 11+9=278\cdot 13+3=212\cdot 17+13$
$=$ $190\cdot 19+7=157\cdot 23+6=124\cdot 29+21=116\cdot 31+21=97\cdot 37+28=88\cdot 41+9=84\cdot 43+5$
$=$ $76\cdot 47+45=68\cdot 53+13=61\cdot 59+18$
@@
$q=$ $2010\cdot 2+1=1340\cdot 3+1=804\cdot 5+1=574\cdot 7+3=365\cdot 11+6=309\cdot 13+4=236\cdot 17+9$
$=$ $211\cdot 19+12=174\cdot 23+19=138\cdot 29+19=129\cdot 31+22=108\cdot 37+25=98\cdot 41+3$
$=$ $93\cdot 43+22=85\cdot 47+26=75\cdot 53+46=68\cdot 59+9=65\cdot 61+56$
@@
$P_{2}^{2}=c_{2}^{2}\mathcal{O}_{K}+\left(2+\sqrt{d}\right)\mathcal{O}_{K}$, $P_{2}^{3}=c_{2}^{3}\mathcal{O}_{K}+\left(27+\sqrt{d}\right)\mathcal{O}_{K}$, $P_{2}^{5}=c_{2}^{5}\mathcal{O}_{K}+\left(1402+\sqrt{d}\right)\mathcal{O}_{K}$,
$P_{3}^{2}=c_{3}^{2}\mathcal{O}_{K}+\left(30+\sqrt{d}\right)\mathcal{O}_{K}$, $P_{4}^{2}=c_{4}^{2}\mathcal{O}_{K}+\left(623+\sqrt{d}\right)\mathcal{O}_{K}$, $P_{6}^{2}=c_{6}^{2}\mathcal{O}_{K}+\left(943+\sqrt{d}\right)\mathcal{O}_{K}$,
$P_{7}^{2}=c_{7}^{2}\mathcal{O}_{K}+\left(1188+\sqrt{d}\right)\mathcal{O}_{K}$, $P_{10}^{2}=c_{10}^{2}\mathcal{O}_{K}+\left(2501+\sqrt{d}\right)\mathcal{O}_{K}$, $P_{15}^{2}=c_{15}^{2}\mathcal{O}_{K}+\left(10898+\sqrt{d}\right)\mathcal{O}_{K}$,
$P_{20}^{2}=c_{20}^{2}\mathcal{O}_{K}+\left(25908+\sqrt{d}\right)\mathcal{O}_{K}$, $P_{21}^{2}=c_{21}^{2}\mathcal{O}_{K}+\left(2211+\sqrt{d}\right)\mathcal{O}_{K}$, $P_{22}^{2}=c_{22}^{2}\mathcal{O}_{K}+\left(22868+\sqrt{d}\right)\mathcal{O}_{K}$,

### v2

# On counterexamples to Mordell’s Pellian Equation Conjecture and the AAC-Conjecture: a non-computer based approach

@@ Abstract.
In this note, we discuss recently discovered counterexamples to Mordell’s Pellian Equation Conjecture and the Ankeny-Artin-Chowla-Conjecture. We provide a verification of the counterexample to Mordell’s Pellian Equation Conjecture that can be checked with marginal computer assistance.
@@ 1. Introduction and main result
We fix notation which is needed to formulate the Ankeny-Artin-Chowla-Conjecture and Mordell’s Pellian Equation Conjecture.
@@
Observe that $\mathcal{O}_{K}=\mathbb{Z}[\omega]=\mathbb{Z}\oplus\omega\mathbb{Z}$. Let $p\in\mathbb{P}$. Then $p$ is called ramified in $\mathcal{O}_{K}$ if $p\mathcal{O}_{K}$ is the square of a maximal ideal of $\mathcal{O}_{K}$. Moreover, we say that $p$ splits in $\mathcal{O}_{K}$ if $p\mathcal{O}_{K}$ is the product of two distinct maximal ideals of $\mathcal{O}_{K}$.
We recall two (long standing) conjectures.
Conjecture 1.1 (Ankeny-Artin-Chowla-Conjecture).
Let $x,y\in\mathbb{N}_{0}$ be such that $\varepsilon=x+y\omega$. If $d\in\mathbb{P}$ and $d\equiv 1\mod 4$, then $d\nmid y$.
Conjecture 1.2 (Mordell’s Pellian Equation Conjecture).
Let $x,y\in\mathbb{N}_{0}$ be such that $\varepsilon=x+y\omega$. If $d\in\mathbb{P}$ and $d\equiv 3\mod 4$, then $d\nmid y$.
We briefly discuss the history and the significance of both conjecture. The Ankeny-Artin-Chowla-Conjecture was first mentioned in 1952 (see [2, page 480]) and it was subsequently investigated in a multitude of papers [1, 3, 7, 8, 11, 13, 19, 21]. Mordell’s Pellian Equation Conjecture was stated by L. J. Mordell in 1961 (see [14, page 283]) and independently by A. A. Kiselev and I. Sh. Slavutskiĭ in 1959 (see [10]). It was investigated in various papers and by many authors (for instance, see [3, 5, 12, 19, 21]).
Both conjectures were recently disproved by counterexamples which were found by a long numerical computer search [17, 18]. For the methods, algorithms and programs that were used to find (and check) these counterexamples, we refer to [17, Section 3] and [18, Section 2]. Note that the algorithms (that were used in [17, 18] to check the counterexamples) require a multitude of computational steps and make (at least to some extent) use of continued fraction expansions. For instance, the fundamental unit of $\mathcal{O}_{K}$ for $d$ in Theorem 1.3 is larger than $10^{1880000}$ and the period length of the continued fraction expansion of $\sqrt{d}$ is $3650856$.
The purpose of this note is to demonstrate how to verify the given counterexamples with manageable effort and only minor computer assistance. We do this in detail for the counterexample to Mordell’s Pellian Equation Conjecture. The verification of the counterexample to the AAC-Conjecture can be done along similar lines. Also note that we do not use continued fraction expansions, and thus we get additional independent verifications of the counterexamples.
Note that both the Ankeny-Artin-Chowla-Conjecture and Mordell’s Pellian Equation Conjecture have ties with some problems in (commutative) algebra. Let $\mathcal{O}$ be an order in $K$. Then $\mathcal{O}$ is said to have torsionfree cancellation if for all torsionfree finitely generated $\mathcal{O}$-modules $A,B,C$ with $A\oplus B\cong A\oplus C$, it follows that $B\cong C$. Moreover, $\mathcal{O}$ is said to have mixed cancellation if for all finitely generated $\mathcal{O}$-modules $A,B,C$ with $A\oplus B\cong A\oplus C$, we have that $B\cong C$. Let $d\in\mathbb{P}$ and let $x,y\in\mathbb{N}_{0}$ be such that $\varepsilon=x+y\omega$. It is well-known that $\mathcal{O}_{d}$ has torsionfree cancellation if and only if $d\nmid y$ [22, Theorem 4.9]. In particular, it follows from [22, Theorem 4.9] and [9, Theorem 5.12] that if $d\geq 7$, then $\mathcal{O}_{d}$ has torsionfree cancellation, but not mixed cancellation if and only $d\nmid y$.
We also want to mention a problem from factorization theory that is connected to the condition $d\mid y$. It was recently studied in [16, 17]. For a profound introduction to factorization theory see [6]. Let $\mathcal{O}$ be an order in $K$. A nonzero nonunit of $\mathcal{O}$ that cannot be written as a product of two nonunits of $\mathcal{O}$ is called an atom of $\mathcal{O}$. It is well-known that every nonzero nonunit of $\mathcal{O}$ is a finite product of atoms of $\mathcal{O}$ (see [6]). Let $x\in\mathcal{O}$ be a nonzero nonunit. If $\ell\in\mathbb{N}$ is such that $x$ is a product of $\ell$ atoms of $\mathcal{O}$, then $\ell$ is called a length of $x$. Let $\mathsf{L}(x)$ be the set of lengths of $x$. Set $L=\min\{r-s\mid y\in\mathcal{O}$ is a nonzero nonunit, $r,s\in\mathsf{L}(y),s<r\}$ (note that we use $\min\emptyset=0$ here for convenience). It is known (see [16]) that $L\leq 2$. We say that $\mathcal{O}$ is unusual if $L=2$. The condition $d\mid y$ determines in many cases whether $K$ has an unusual order (see [17, Proposition 2.4 and Theorem 2.5]).
For a survey on the Conjectures 1.1 and 1.2 and further discussion of their importance, we refer to [20]. For a more detailed study of the property $d\mid y$ for arbitrary squarefree $d\in\mathbb{N}_{\geq 2}$, see [17]. Next, we present the main result of this note.
@@
The proof of Theorem 1.3 in Section 3 was created with computer assistance (and with the help of the programs Mathematica 12.0.0 and Pari/GP 2.15.2). We used the command bnfunits(bnfinit(X2-d,1)) in Pari/GP (together with some elementary algebraic manipulations) to obtain the values in the table (with the entries $a_{i}$, $b_{i}$, $c_{i}$, $d_{i}$, $e_{i}$ and $f_{i}$) below.
Observe that (very dedicated) arithmeticians may be able to verify this proof without computer assistance. We define an arithmetical operation as hard if it is a multiplication of integers $a$ and $b$ with $|a|\geq 10^{3}$, $|b|\geq 10^{3}$ and $|ab|\geq 10^{10}$. All other arithmetical operations in this note (additions, subtractions and multiplications of integers) are defined to be easy. We define additions and subtractions where (at least) one of the members is a single digit integer as trivial. A rough estimate yields that there are approximately 550 hard operations and about 2200 (nontrivial) easy operations in the proof of Theorem 1.3 below. If we assume that (the verification of) a hard operation takes about 15 minutes (on average) and (the verification of) an easy (but nontrivial) operation takes about 3 minutes (on average), then a verification of the proof below is doable in about 250 hours (without computers). If all arithmetical operations are done with computers, then it should only take a modest amount of time (to check all the details).
@@ 2. Preliminaries
We recall a variant of a result of H. C. Pocklington (cf. [4, Theorem 4] and [15]) that will be used to show that $d$ is prime. For the readers’ convenience we include a proof. This proof goes along the lines of the proof of [4, Theorem 4 and Corollary 1].
@@
Let $d=39028039587479$ and let $x,y\in\mathbb{N}_{0}$ be such that $\varepsilon=x+y\omega$. We need to show that $d\equiv 3\mod 4$, $d\in\mathbb{P}$ and $d\mid y$. Since $79=19\cdot 4+3$, we have that $d\equiv 79\equiv 3\mod 4$. Observe that $\omega=\sqrt{d}$, $x,y\in\mathbb{N}$ and $\varepsilon=x+y\sqrt{d}$.
@@
$p=$ $1808\cdot 2+1=1205\cdot 3+2=723\cdot 5+2=516\cdot 7+5=328\cdot 11+9=278\cdot 13+3=212\cdot 17+13=190\cdot 19+7=157\cdot 23+6$
$=$ $124\cdot 29+21=116\cdot 31+21=97\cdot 37+28=88\cdot 41+9=84\cdot 43+5=76\cdot 47+45=68\cdot 53+13=61\cdot 59+18$
@@
$q=$ $2010\cdot 2+1=1340\cdot 3+1=804\cdot 5+1=574\cdot 7+3=365\cdot 11+6=309\cdot 13+4=236\cdot 17+9=211\cdot 19+12=174\cdot 23+19$
$=$ $138\cdot 29+19=129\cdot 31+22=108\cdot 37+25=98\cdot 41+3=93\cdot 43+22=85\cdot 47+26=75\cdot 53+46=68\cdot 59+9=65\cdot 61+56$
@@
$P_{2}^{2}=c_{2}^{2}\mathcal{O}_{K}+\left(2+\sqrt{d}\right)\mathcal{O}_{K}$, $P_{2}^{3}=c_{2}^{3}\mathcal{O}_{K}+\left(27+\sqrt{d}\right)\mathcal{O}_{K}$, $P_{2}^{5}=c_{2}^{5}\mathcal{O}_{K}+\left(1402+\sqrt{d}\right)\mathcal{O}_{K}$, $P_{3}^{2}=c_{3}^{2}\mathcal{O}_{K}+\left(30+\sqrt{d}\right)\mathcal{O}_{K}$,
$P_{4}^{2}=c_{4}^{2}\mathcal{O}_{K}+\left(623+\sqrt{d}\right)\mathcal{O}_{K}$, $P_{6}^{2}=c_{6}^{2}\mathcal{O}_{K}+\left(943+\sqrt{d}\right)\mathcal{O}_{K}$, $P_{7}^{2}=c_{7}^{2}\mathcal{O}_{K}+\left(1188+\sqrt{d}\right)\mathcal{O}_{K}$, $P_{10}^{2}=c_{10}^{2}\mathcal{O}_{K}+\left(2501+\sqrt{d}\right)\mathcal{O}_{K}$,
$P_{15}^{2}=c_{15}^{2}\mathcal{O}_{K}+\left(10898+\sqrt{d}\right)\mathcal{O}_{K}$, $P_{20}^{2}=c_{20}^{2}\mathcal{O}_{K}+\left(25908+\sqrt{d}\right)\mathcal{O}_{K}$, $P_{21}^{2}=c_{21}^{2}\mathcal{O}_{K}+\left(2211+\sqrt{d}\right)\mathcal{O}_{K}$, $P_{22}^{2}=c_{22}^{2}\mathcal{O}_{K}+\left(22868+\sqrt{d}\right)\mathcal{O}_{K}$,